OpenStudy (anonymous):

Could somebody help me with x^2 + 10x + 7 is greater than or equal to 0? I've gotten as far as -5 plus or minus three times the square root of 2.

OpenStudy (compassionate):

This is a quadratic equation. It can be solved with binomials

OpenStudy (alekos):

Are you saying that you got x = -5 + 3sqrt2 & x = -5 - 3sqrt2 ?

OpenStudy (anonymous):

Suppose you would have \[x^2 + 10x + 7 = 0 \] How would you solve it?

OpenStudy (campbell_st):

a simple thing to do is graph toe curve and see where the curve is on or above the x-axis... this will give the inequality

OpenStudy (alekos):

Yes, that would be an easier way. Go to Desmos and graph it up. I think you can even use > symbol in Desmos